In tomatoes, red fruit is dominant to yellow fruit and hairy stems are dominant to hairless (smooth) stems. The two genes controlling these traits are inherited in normal Mendelian fashion and are not linked.

 

a)   Produce a suitable key for this information in terms of the individual alleles of the genes given.                                                                                 

 

 

b)   What are the possible genotypes of plants showing the following forms:    

                                                                                                           

 

i)              Red fruit and smooth stems

 

ii)             Yellow fruit and hairy stems

 

iii)           Red fruit and hairy stems

 

c)   On crossing a plant with red fruit and hairy stem, with one that was yellow fruit and hairy stem, the following progeny were obtained:

 

           30 red, hairy / 11 red, smooth / 29 yellow, hairy / 9 yellow, smooth

 

Give a reasoned explanation along with the use of full genetic diagrams to proof your answer, to explain these results in terms of the simultaneous transmission of characteristics.                         


1
Expert's answer
2021-02-15T03:18:36-0500

a) Red = R ; Yellow= r

Hairy stems = H ; hairless stems = h



b)

Pr(Red fruit) = 3/4

Pr(Yellow fruit) = 1/4

Pr(Hairy stems) = 3/4

Pr(Smooth stems) = 1/4


(i) Red stems × smooth stem

3/4 × 1/4 = 3/16


(ii) Yellow fruit × Hairy stem

1/4 × 3/4 = 3/16


(iii) Red fruit × Hairy stem

3/4 × 3/4 = 3/16


c) After drawing a Punnet square and inserting the appropriate values, compare with the values in the question.